OpenStudy (anonymous):

A girl is now one-third as old as her mother. In three years, she will be two-fifths as old as her mother will be. What are their present ages? girl is 9; mom is 27 girl is 18; mom is 54 girl is 25; mom is 75

OpenStudy (anonymous):

help

OpenStudy (aum):

Let G be the current age of the girl and M be the current age of the mother. "A girl is now one-third as old as her mother." G = 1/3 * M ---- (1) In three years, the girl will be G+3 years old and mother will be M+3 years old. " In three years, she will be two-fifths as old as her mother will be." G+3 = 2/5 * (M+3) ----- (2) Put G = M/3 from (1) into (2) and solve for M. Then solve for G.

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

a b or c

OpenStudy (aum):

Can't give out answers. Only can help you learn.